could someone please help me with this equation????
use elimination to solve:
6x-y/4=-9
and 3x+2y=-30

Answer by joey_quizan(15)

given:
First equation : 6x -y/4 = -9
Second Equation: 3x + 2y = -30

Solution:
Step 1: multiply first equation by 8 to eliminate y
8[6x - y/4 = -9]

48x - 2y = -72 --- eq 3
Step 2: Add eq 3 and second equation
48x - 2y = -72 --- eq 3
+ 3x + 2y = -30 --- second equation
-------------------
51x + 0 = -102

51x = -102
--- ----
51 51

x = -2
Step 3: substitute the value of x to second equation to find the value of y

3x + 2y = -30 --- second equation
3(-2) + 2y = -30
-6 + 2y = -30 --- transpose -6 to the other side of the equation
2y = -30 + 6
2y = -24
-- ---
2 2
y = -12
Check if the answere is correct substitute the value of x and y to first equation
6x - y/4 = -9 --- first equation
6(-2) - (-12/4) = -9
-12 + 3 = -9
-9 = -9 --- therefore the answere is corre
Answere:

x = -2 ; y = -12